You and three friends stand at the corners of a square whose sides are 8.0 m long in the middle of the gym floor, as shown below. You take your physics book and push it from one person to the other. The book has a mass of 1.5 kg, and the coefficient of kinetic friction between the book and the floor is μ_k=0.25

a. The book slides from you to Beth and then from Beth to Carlos, along the lines connecting these people. What is the work done by friction during this displacement?

b. You slide the book from you to Carlos along the diagonal of the square. What is the work done by friction during this displacement?

c. You slide the book to Kim, who then slides it back to you. What is the total work done by friction during this motion of the book?

d. Is the friction force on the book conservative or non-conservative? Explain.



1
Expert's answer
2021-03-10T17:19:35-0500

a)


"W=3\mu mgd\\W=3(0.25)(9.8)(8)(1.5)=88.2\ J"

b)


"W=\sqrt{2}\mu mgd\\W=\sqrt{2}(0.25)(9.8)(8)(1.5)=41.6\ J"

c)


"W=2\mu mgd\\W=2(0.25)(9.8)(8)(1.5)=58.8\ J"

d) A nonconservative force is one for which work depends on the path taken. Friction is a good example of a nonconservative force. Friction, for example, creates thermal energy that dissipates, removing energy from the system.